[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Jul 26, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Wed Jul 26 13:15:12 UTC 2017


Here is today's summary of NHC computer operations:
--- PS

----------

TAFB:
1.) 1535 L EPAC 12Z NWPS run started 1419Z, finished 1719Z but no 12Z NWPS  
grids. Matt Sardi/Chris Mello investigating. This was resolved yesterday as  
Mello was in the middle of reconfiguring EPAC domain. He is however  
following up problems reported overnight. See below.

2.) NWPS for Pacific took the length of time sent in JL email, although it  
took an additional 15 minutes or so following WRKNWP completion message to  
show in GFE. The 00z Pacific run was randomly missing timestep hours 24,  
27, 30, 33, 36, 39, 42, 132, 147, 150, 159, 162, 165, 168. The 06z Pacific  
run came in complete. Chris Mello is investigating this.
HSU:
1.) JTWC called and said they were having trouble getting the forecasts for  
the EPac storms into their ATCF. Monica and Matt Sardi notified.

2.) GOES16 imagery stopped coming in after 1441Z until 1505Z-- then it all  
came in at once. Talked to Mello and he coordinated with Monica to fix  
problem.

3.) FSSE continues to come late. Lengthy email exchanged with responsible  
parties on this yesterday. Working on a resolution. Craig working with Mike  
on this.

Other items:
4.) Brian Maher installed the libraries needed to migrate Anu's Flight  
Tracking software to Compuserv. But there might be some coding issues that  
need to be worked out.

Craig, work with Anu and Brian to get this resolved. After speaking with  
Brian this morning it is clear this is the next step needed to see if we  
can get to the the bottom of it.
NHC RFCs:
1. ) Work on NetApp Migration continues this week.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170726/ebe6bf06/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list